solve for p

UTHEN Question 6 For 6-9 Given: q = $1,000 Is = $290 1x = $1,490 E(1) = $1,280 6. Solve for p 12.8% 14.9% 2.9% 17.5%
Answer #1

Answer:

Given:

q = $1,000

Is = $290

Ih = $1,490

E(I) = $1,280

Expected income, E(I) = Is * probability of being sick + Ih * probability of being healthy

Of course, probability of being healthy = 1 - probability of being sick

E(I) = 1280 = 290*P + 1490*(1 - P)

1280 = 290*P + 1490 - 1490*P

(1490 - 290)*P = 1490 - 1280

1200*P = 210

So, P = 210/1200 = 0.175

P = 17.5%

Hence Correct option is 17.5%